Home‎ > ‎Reports‎ > ‎Crew099‎ > ‎

day13


Crew 99 Log Book for February 10, 2011

Photos of the day





Commander's Check-In Report

Donna Viola Reporting

Crew Physical Status: Good

Time Departed/Returned from EVA:

EVA #23, 10:00-12:00;

EVA #24, 14:00-16:00

Brief Narrative of Field Mission Results: Collecting geological data,

soil sample collection, endolith sample collection

EVA Data/Interpretations: As per EVA reports

Engineering/Hab Maintenance: Nothing major to report.

Report Transmission Schedule: Engineering, EVA #23, EVA #24,

Commander's. Engineering report will probably be late, since both

engineers have been working on the balloon launch since ~5pm.

Plans for Tomorrow: To be decided

Inventory: None

Support Requested: None

Miscellaneous: Currently working on tethered balloon launch


Commander's Report

Donna Viola Reporting

Today, we planned on wrapping up our science for the rotation. This morning, Heidi, Andie, and Max went to finish gathering geological data, and to collect a soil sample that we plan on testing for life with our LAMBDA project after our rotation.

Andie and Heidi then prepared some Velveeta macaroni and cheese for lunch - which gave us plenty of fuel for the afternoon EVA! Sukrit, Kevin, Heidi, and myself went up north to collect our last soil sample, as well as my last endolith samples. We also admired gypsum deposits along the way - much ''geologizing'' was done, as we all returned to the Hab with pockets full of rocks to bring home with us.

When we got back, Max was busy finishing up his preparations for his tethered balloon launch. Since the winds were not cooperating, he needed to scrub his plans for an actual launch at the last minute, and redefine the project as a tethered launch.

 At the moment, Max and Kevin are out working with the balloon, while Sukrit is trying to use the observatory to track it. Updates on this will be provided as they become available.


Engineering Report

Kevin Newman Reporting

*Generator/Electricity:*

·          Kitty  (Diesel Generator) (Used? Yes/No): Yes

·          Honey   (Gas Generator) (Used? Yes/No): No

·          Battery Charge Level: Checked by DG

o    10.59 V everstart

o    12.37 V armor plate

o    13.32 V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good


*Fuel Status:*

·          Diesel (Full/0.75/0.5/0.25/Empty): 0.90

·          Propane Tank (Full/0.75/0.5/0.25/Empty): 0.68

·          Gasoline Tank (Full/0.75/0.5/0.25/Empty): 14 Gallons

 Notes/Comments:


*Oil:*

·          Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

·          Honey Oil Quantity (# of Quarts in storage): Leave it to DG

·          ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

·          Car Oil Quantity (# of Quarts in storage):


*Water Status:*

·           Outside Potable Water Tank Level (inches from bottom): 14

·           Trailer Potable Water Tank Level (inches from bottom): 5/6 Full

·           Hab Potable Water Tank Level (inches from bottom): 7

·           Potable Water Meter Reading: 418962

Notes/Comments:


GreenHab:

·          Condition of plants in Tank 1: Nothing

·          Condition of plants in Tank 2: Ok

·          Condition of plants in Tank 3: Good

·          Greenhab notes/ comments:


Observatory:

·          Telescope Used? (Yes/no) Yes

 Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

·          Used (yes/no): No

·          Oil Checked (yes/no): No

·          Tire Status: Not checked

·          Odometer at end of day: 123611.6

·          Notes/Comments on Hab Car:


ATV 1 (Opportunity):

·          Used: Yes

·           Oil Checked: Yes

·           Fuel Consumed: Replaced

·           Tire Status: Filled

·           Comments:


 ATV 2 (Spirit):

·          Used: Yes

·          Oil Checked: Yes

·          Fuel Consumed: Replaced

·          Tire Status: Filled

·          Comments:


ATV 3 (Viking I):

·           Used: Yes

·           Oil Checked: Yes

·           Fuel Consumed: Replaced

·           Tire Status: Filled

·           Comments: Difficult to start the first time, especially in sim,

but gets easier as the engine warms up. The operator of this ATV must know how to use the choke and the pull cord to start it every time.


ATV 4 (Viking II):

·           Used: Yes

·           Oil Checked: Yes

·           Fuel Consumed: Replaced

·           Tire Status: Filled

·           Comments:


Heating and Ventilation:

·          Thermostat setting upstairs (Farenheit): 70

·          Actual temperature upstairs (Farenheit): 74

·          Thermostat setting downstairs (Farenheit): 65

·          Actual temperature downstairs (Farenheit): 64


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: Behaving well


General Engineering

1. Notes/Comments:

I tried to send this report a few times last night, each time getting the

following message from gmail:

We're writing to let you know that the group you tried to contact (mdrs-mission-support) may not exist, or you may not have permission to post messages to the group. A few more details on why you weren't able to post:

 * You might have spelled or formatted the group name incorrectly.

 * The owner of the group may have removed this group.

 * You may need to join the group before receiving permission to post.

 * This group may not be open to posting.


EVA 23 Report

Heidi Beemer Reporting

Time: 1000-1200

Crew Members: Heidi Beemer, Andie Gilkey

Site Location:  0519731   4251449 , 0520249  4251141 and 0518156 4250273

Transit Mode: ATV

Objectives:

Collect RMR data for location shelter site two.

Take photos at all three sites.

Collect soil sample for MFC. 

Results: Photos for use of sterophotography were taken at all three sites. At the second site discontinuities were measured for strike and dip as well as distance apart from each other.

Soil was collected for LAMBDA at 0520160  4251186

Results to follow in Geology Report

Lessons Learned: None


EVA 24 Report

Donna Viola Reporting

Time 14:00-16:00

Crew Members: Heidi Beemer, Donna Viola, Sukrit Ranjan, Kevin Newman

Site Location:  517884, 4254729; 518802, 4253236

Transit Mode: ATVs, foot

Objectives: Collect soil sample for microbial fuel cell; collect endolith specimens

Results: Successfully collected appropriate samples; will be analyzed back at our home institutions

Lessons Learned:  Science is fun!


Science Report (Geology)

Heidi Beemer Reporting

EVA#: 22, 23

Crew Members Involved: Heidi Beemer

Site 1 Location: 0518156  4250273

Site 2 Location: 0519731  4251449

Site 3 Location: 0520249  4251141

Transit Mode: ATV

Objectives: Calculate geomechanics of three sites using  RMR (Bieniawski, 1972). These parameters will tell the engineer the behavior of the rock, providing quantitative data for engineering design.

Due to the unique geology of “Mars” in the San Rafael Swell, there are many locations in which tunnel type shelters can be built for expansion during colonization. The purpose of this investigation is to calculate a RMR rating of good rock or better (61-100) for three possible shelter locations.

Results: A RMR rating of 63 was determined from the first site. This was determined on the following parameters:

Uniaxial compressive strength of rock material: Medium High 50-100 MPa................................7

 Rock Quality Designation (RQD): Fair Quality 70%.................................................13                      

Spacing of discontinuities: Moderate 200-600mm.................................................………………………………………10

Condition of discontinuities: Moderately weathered…………………………………….................................................20

Groundwater conditions: Completely dry.................................................15

Orientation of discontinuities: Drive w/ dip Favorable…………………………………-2

                                                                                                                        Total: 63

This rating is a class II with a description of good rock.

A RMR rating of 74 was determined from the second site. This was determined on the following parameters:

Uniaxial compressive strength of rock material: Moderate  25-50 MPa..............................4

 Rock Quality Designation (RQD): Excellent 100%..........................................20                        

Spacing of discontinuities: Moderate 0.6-2m……………………….................................................15

Condition of discontinuities: Moderately weathered……………………………….................................................20

Groundwater conditions: Completely dry.................................................15

Orientation of discontinuities: Drive w/ dip Very Favorable………………………0

                                                                                                                        Total: 74

This rating is a class II with a description of good rock.

A RMR rating of 65 was determined from the third site. This was determined on the following parameters:

Uniaxial compressive strength of rock material: Medium High 50-100 MPa……….............................7

 Rock Quality Designation (RQD): Fair Quality 70%.................................................13                       

Spacing of discontinuities: Moderate 200-600mm.................................................………………………………………10

Condition of discontinuities: Slightly weathered………………………………………….................................................25

Groundwater conditions: Completely dry.................................................15

Orientation of discontinuities: Drive w/ dip Unfavoraable……………………………….................................................10

                                                                                                                        Total: 65

This rating is a class II with a description of good rock.


Journalist Report

Sukrit Ranjan Reporting

We were thoughtful today as we completed our research projects. After EVAs to collect more soil and rock samples and an abortive attempt to get a geocache (alas! for we may not go on the highway), we suited up for the balloon experiment.

For a little while, we got to relive our childhoods, towing a balloon that was much large than us and carrying an instrument payload like children with a carnival balloon. Good data was collected from the Musk Observatory.

After 7 hours of being outside in the freezing cold, it was  a relief to get indoors. It is still sinking in that this was our penultimate day on Mars. It can't be over so quickly, can it?